OpenStudy (anonymous):

What is the solution to p over zero point three - 2 = 6?

OpenStudy (anthonyym):

Do you mean \[\frac{ p }{ 0.3 }-2=6\]

OpenStudy (texaschic101):

(p/0.3) - 2 = 6 p/0.3 = 6 + 2 p/0.3 = 8 p = 0.3 * 8 p = 2.4

OpenStudy (anthonyym):

[Option 2] Or do you mean \[\frac{ p }{ 0.3 - 2 }=6\]

OpenStudy (texaschic101):

true...it depends on how the problem is written

OpenStudy (anonymous):

option 1

OpenStudy (texaschic101):

then the way i did it gets u....p = 2.4

OpenStudy (anonymous):

thank you texas chick the way you put it made me realise ive done this before. lol

OpenStudy (texaschic101):

yw